# MASt3R-Fusion: Integrating Feed-Forward Visual Model with IMU, GNSS for High-Functionality SLAM

### v2

@@ VII-A KITTI-360 Dataset
We first focus on the performance of data association. Fig. 8 shows the performance of temporal tracking and cross-temporal matching based on the feed-forward model. For temporal tracking, it can be observed that the feed-forward model provides dense pixel-level associations and is able to directly establish reliable correspondences between frames with large intervals. Compared with traditional multi-frame feature tracking methods, such long-span associations provide more direct and stronger constraints. Besides, moving objects (e.g., cars) can be effectively excluded in the association based on 3D awareness. Regarding cross-temporal matching (for loop closure), the 3D priors embedded in the feed-forward model makes dense matching possible even under extreme viewpoint differences. To be specific, it enables imaginative data associations in cases where the viewpoint difference exceeds 90°, or even when the views are completely opposite—something that is difficult to achieve with conventional feature-based matching methods.
The focus is then put on pose estimation. Considering the relatively large scale, we evaluate the odometry performance using relative pose metrics, following the KITTI[40] odometry benchmark. The results are listed in TABLE I. The corresponding representative trajectory of the data sequence is shown in Fig. 9(a). It can be seen that the proposed MASt3R-Fusion achieves accurate pose estimation, yielding lower relative translation errors compared with both traditional indirect/direct VIO methods and learning-based VIO approaches. This improvement can be attributed to the powerful data association enabled by the feed-forward model, as well as the effective pointmap-alignment-based visual constraints. In addition, the visual-only MASt3R-SLAM almost fails to complete SLAM at such a large scale, whose scale estimation is severely affected by the imperfect information of pointmap regression. The incorporation of inertial information effectively mitigates these errors and fully leverages the strong dense visual constraints for stable metric-scale pose estimation.
@@
ORB-SLAM3
MASt3R-Fusion
@@
Afterwards, we evaluate the global SLAM performance, by activating loop closure and global optimization. For comparison, we test ORB-SLAM3 which supports loop closure and BA-based global optimization. We simply account the absolute translation errors (ATEs) for evaluation[41], which are listed in TABLE II. The corresponding trajectory is shown in Fig. 9(b). It can be observed that the proposed MASt3R-Fusion achieves a significantly lower ATE than ORB-SLAM3. This is partly due to the higher odometry accuracy, and partly because the proposed method incorporates stronger and richer loop closure information. As illustrated in Fig. 8, loop closures under large viewpoint differences allow the system to maintain pose consistency even when driving in opposite directions, which provides favorable conditions for globally consistent mapping.
One of the functionalities of the proposed system is real-time 3D spatial percpetion with metric scale awareness. In Fig. 10, we qualitatively evaluate this by comparing the proposed system with learning-based dense VIO[30] system and metric depth inference model[42]. From the results, it can be observed that all the three methods are capable of achieving dense perception of the environment with mertic scale. Among them, DBA-Fusion primarily relies on BA for reconstructing the environment structure but lacks 3D priors, which leads to scattered points and poor handling of dynamic objects in the scene. The monocular depth model Metric3D v2, through specialized training, demonstrates strong depth estimation performance, particularly in terms of surface normals and edges. However, since it fully depends on inference, objects with ambiguous depth cues can exhibit significant depth/scale errors. In contrast, MASt3R-Fusion preserves the overall 3D structure while leveraging binocular geometry, thereby attaining relatively stable and complete real-time 3D perception performance.
@@ VII-B SubT-MRS Dataset
Similarly, we first focus on the performance of data association. It can be seen from Fig. 11 that in unconventional scenarios (such as caves), the feed-forward model still provides reliable dense feature associations, even when the scene texture is only weakly distinguishable. The matching performance under large viewpoint differences remains excellent, even surpassing the range discernible by the human eye, thereby offering strong potential constraints for global SLAM.
Afterwards, we conducted tests on these data sequences for both real-time and global V-I SLAM. The results of the pose estimation are presented in TABLE III and TABLE IV.
From the real-time SLAM results, it can be observed that MASt3R-Fusion achieves significantly higher pose estimation accuracy than existing methods, providing stable metric-scale odometry performance. In terms of global SLAM performance, the aggressive loop closure detection delivers impressive results, leading to a substantial reduction in ATEs. This part of the experiments largely validates the generalizability of the feed-forward model approach on open scenarios when combined with IMU information.
@@
ORB-SLAM3
MASt3R-Fusion
@@
VII-C Wuhan Urban Dataset
In addition to the evaluation of the V-I system, to further evaluate the algorithm’s performance in real-world road scenarios (with dynamic objects) and to consider its integration with global GNSS measurements, we turn to the self-collected multi-sensor dataset acquired in Wuhan City. The dataset contains two sequences, with their trajectories shown in Fig. 13. We take the classical VINS-Fusion as the baseline, whichs supports both real-time pose estimation and global fusion with GNSS.
@@ VII-C Wuhan Urban Dataset
First, we tested several different pose estimation schemes based on V-I data, including real-time VIO and global SLAM. The corresponding trajectory is shown in Fig. 14. It can be observed that in large-scale road environments (with long straight segments), the traditional VIO algorithm suffers from significant scale estimation errors due to weak observability[43]. In contrast, the proposed MASt3R-Fusion method, by combining probabilistic marginalization with Sim(3)-based visual constraints, can effectively control and correct scale drift. Besides, after introducing loop closure constraints, the position error caused by attitude drift is greatly reduced, and the trajectory accuracy is further improved.
Second, the global positioning performance with GNSS integration is focused on. The positioning errors are depicted in Fig. 15. From the results, it can be observed that GNSS RTK exhibits many gross errors and periods of unavailability, especially in sequence (a). For VINS-Fusion, in which GNSS positions are fused with VIO-provided relative poses loosely through a pose graph, the position estimation errors during GNSS-degraded periods cannot be effectively suppressed, since the position residuals cannot be probabilistically distributed across all poses. This further undermines the detection and mitigation of GNSS gross errors. In contrast, in the proposed MASt3R-Fusion, the complete and precise V-I information is fully preserved into the factor graph, effectively resisting GNSS gross errors through iterative optimization and maintaining decimeter-level accuracy.
Moreover, to enrich the variability of GNSS observation conditions in the tests, we use simulated GNSS data based on ground truth with intermittent outages to assess the robustness of the algorithm. The results are shown in Fig. 16, which further demonstrate the effectiveness of the proposed method in recovering globally optimal trajectories. Impressively, with 100-second GNSS outages, the global factor graph can still achieve mostly sub-meter level trajectory smoothing, and the introduction of cross-temporal associations further reduces the overall error.
VII-D Reconstruction Results
@@
Figure 15: Horizontal errors of different absolute positioning schemes with GNSS RTK integration. The bolded numbers indicate the root mean square errors (RMSEs).
Figure 16: Horizontal errors of different absolute positioning schemes with simulated GNSS integration. The gray crosses indicate simulated GNSS outages (100-second). The bolded numbers indicate the RMSEs.
VIII Conclusion
@@ Acknowledgments
The implemented MASt3R-Fusion is developed by the GREAT Group, School of Geodesy and Geomatics, Wuhan University. The numerical calculations in this paper have been done on the supercomputing system at the Supercomputing Center of Wuhan University.

### v3

@@ VI-A KITTI-360 Dataset
We first focus on the performance of data association. Fig. 8 shows the performance of temporal tracking and cross-temporal matching based on the feed-forward model. For temporal tracking, it can be observed that the feed-forward model provides dense pixel-level associations and is able to directly establish reliable correspondences across frames with large intervals. Compared with traditional multi-frame feature tracking methods, such long-span associations provide directer and stronger constraints. Besides, moving objects (e.g., cars) can be effectively excluded in the association based on 3D awareness. Regarding cross-temporal matching (for loop closure), the 3D priors embedded in the feed-forward model makes dense matching possible even under extreme viewpoint differences. To be specific, it enables imaginative data associations in cases where the viewpoint difference exceeds 90°, or even when the views are completely opposite—something that is difficult to achieve with conventional feature-based matching methods.
The focus is then put on pose estimation. Considering the relatively large scale, we evaluate the odometry performance using relative pose metrics, following the KITTI[42] odometry benchmark. The results are listed in Table I. The corresponding trajectories of representative data sequences are shown in Fig. 9(a). It can be seen that the proposed MASt3R-Fusion achieves accurate pose estimation, yielding lower relative translation errors (RTEs) compared with both traditional indirect/direct VIO methods and learning-based VIO approaches. The average RTE is 43.0% lower than DM-VIO and 17.7% lower than DBA-Fusion, which shows stable performance even for weak-excitation sequences (e.g., 0003 and 0010). This improvement can be attributed to the powerful data association enabled by the feed-forward model, as well as the effective pointmap-alignment-based visual constraints. In addition, the visual-only MASt3R-SLAM almost fails to complete SLAM at such a large scale, whose scale estimation is severely affected by the imperfect information of pointmap regression. The incorporation of inertial information effectively mitigates these errors and fully leverages the strong dense visual constraints for stable metric-scale pose estimation.
@@
VGGT
-Long*
ORB
-SLAM3
MASt3R
-Fusion
@@
    Visual-only, scaled using Sim(3)-based global alignment[43].
Afterwards, we evaluate the global SLAM performance by activating loop closure and global optimization. For comparison, we test ORB-SLAM3 which supports loop closure and BA-based global optimization, and VGGT-Long[44] which utilizes state-of-the-art feed forward model[4] and chunked processing strategy for long-sequence reconstruction. We simply account the absolute translation errors (ATEs) for evaluation[43], as listed in Table II. The corresponding trajectory is shown in Fig. 9(b). It can be observed that the proposed MASt3R-Fusion achieves a significantly lower ATE than ORB-SLAM3. The average ATE/length achieves 0.05%, which is quite impressive considering that only monocular V-I data is used. This is partly due to the higher odometry accuracy, and partly because the proposed method incorporates stronger and richer loop closure information. As illustrated in Fig. 8, loop closures under large viewpoint differences allow the system to maintain pose consistency even when driving in opposite directions, which provides favorable conditions for globally consistent mapping. As for the visual-only VGGT-Long, although the multi-view feed-forward model is employed, its effectiveness is somewhat limited. The significant scale drift of the visual system in large-scale scenes makes the global optimization challenging, and makes it sensitive to false loop closures, yielding to conservative strategies. In contrast, the incorporation of IMU inforamtion in MASt3R-Fusion allows the potential of feed-forward visual models to be fully realized.
One of the functionalities of the proposed system is real-time 3D spatial percpetion with metric scale awareness. In Fig. 10, we qualitatively evaluate this by comparing the proposed system with learning-based dense VIO[34] system and metric depth inference model[45]. From the results, it can be observed that all the three methods are capable of achieving dense perception of the environment with mertic scale. Among them, DBA-Fusion primarily relies on BA for reconstructing the environment structure but lacks 3D priors, which leads to scattered points and poor handling of dynamic objects in the scene. The monocular depth model Metric3D v2 (ViT-Large), through specialized training, demonstrates strong depth estimation performance, particularly in terms of surface normals and edges. However, since it fully depends on inference, objects with ambiguous depth cues can exhibit significant depth/scale errors. In contrast, MASt3R-Fusion preserves the overall 3D structure while leveraging binocular geometry, thereby attaining relatively stable and complete real-time 3D perception performance.
@@ VI-B SubT-MRS Dataset
Similarly, we first focus on the performance of data association. It can be seen from Fig. 11 that in unconventional scenarios (such as caves), the feed-forward model still provides reliable dense pixel associations, even when the scene texture is only weakly distinguishable. The matching performance under large viewpoint differences remains excellent, even surpassing the range discernible by the human eye, thereby offering strong potential constraints for global SLAM.
Afterwards, we conducted tests on these data sequences for both real-time and global V-I SLAM. The results of the pose estimation are presented in Table III and Table IV, and corresponding trajectories are shown in Fig. 12.
@@
VGGT
-Long*
ORB
-SLAM3
MASt3R
-Fusion
@@
    Visual-only.
From the real-time SLAM results, it can be observed that MASt3R-Fusion achieves significantly higher pose estimation accuracy than existing methods, providing stable metric-scale odometry performance. In terms of global SLAM performance, the aggressive loop closure detection delivers impressive results, leading to a substantial reduction in ATEs. This part of the experiments largely validates the generalizability of the feed-forward model approach on open scenarios when incorporated with IMU information.
VI-C Wuhan Urban Dataset
In addition to the evaluation of the V-I system, to further evaluate the algorithm’s performance in real-world road scenarios (with more dynamic objects) and to consider its integration with global GNSS measurements, we turn to the self-collected multi-sensor dataset acquired in Wuhan City. The dataset contains two sequences, with their trajectories shown in Fig. 13. The experimental vehicle is equipped with an RGB camera, an ADIS16470 IMU, a Septentrio AsterRx4 GNSS receiver and a tactical-grade IMU for ground-truth trajectory generation. A nearby GNSS base station is available for differential GNSS (DGNSS) processing. We take the classical VINS-Fusion as the baseline, whichs supports both real-time pose estimation and global fusion with GNSS. We also test DBA-Fusion which employs a learning-based visual frontend and supports real-time V-I(/GNSS) integrated navigation.
@@ VI-C Wuhan Urban Dataset
First, we tested several different pose estimation schemes based on V-I data, including real-time VIO and global SLAM. The corresponding trajectories are shown in Fig. 14, with the relative pose estimation errors listed in Table V. It can be observed that in large-scale road environments (with long straight segments), the traditional VIO algorithm suffers from significant scale estimation errors due to weak observability[46]. This problem remains for DBA-Fusion, although the dense visual information does improve the VIO performance to some extent. In contrast, the proposed MASt3R-Fusion method, by combining probabilistic marginalization with Sim(3)-based visual alignment constraints, can effectively control and correct the scale drift. Besides, after introducing loop closure constraints, the position error caused by attitude drift is greatly reduced, and the trajectory accuracy is further improved.
TABLE V: Relative Pose Errors of Different VIO Schemes on KITTI-360 Dataset.
Seq.
VINS
-Fusion
DBA
-Fusion
M-Fus.
w/o loop
M-Fus.
w/ loop
$t_{rel}$1
$r_{rel}$1
$t_{rel}$
$r_{rel}$
$t_{rel}$
$r_{rel}$
$t_{rel}$
$r_{rel}$
2.07
0.11
1.45
0.11
1.20
0.11
0.53
0.07
5.85
0.13
3.01
0.13
1.68
0.13
0.92
0.06
  • 1
    $t_{rel}$ in %, $r_{rel}$ in /100 m.
Second, the global positioning performance with GNSS integration is focused on. The positioning errors are depicted in Fig. 15 and listed in Table VI. From the results, it can be observed that GNSS RTK exhibits many gross errors and periods of unavailability, especially in sequence (a). For VINS-Fusion, in which GNSS positions are fused with VIO-provided relative poses loosely through a global pose graph, the position estimation errors during GNSS-degraded periods cannot be effectively suppressed, since the position residuals cannot be probabilistically distributed across all poses. This further undermines the detection and mitigation of GNSS gross errors. For DBA-Fusion, although the GNSS inforamtion can assist in the V-I subsystem, it only supports real-time fusion and cannot provide global optimal solutions. In contrast, in the proposed MASt3R-Fusion, the complete and precise V-I information is fully preserved into the factor graph, effectively resisting GNSS gross errors through iterative optimization and maintaining decimeter-level accuracy through the period.
Moreover, to enrich the variability of GNSS observation conditions in the tests, we use simulated GNSS data based on ground truth with intermittent outages to assess the robustness of the algorithm. The results are shown in Fig. 16 and listed in Table VI, which further demonstrate the effectiveness of the proposed method in recovering globally optimal trajectories. Impressively, under 100-second GNSS outages, the global factor graph can still achieve mostly sub-meter level trajectory smoothing, and the introduction of cross-temporal associations further reduces the overall error.
Figure 15: Horizontal errors of different absolute positioning schemes with GNSS RTK integration.
Figure 16: Horizontal errors of different absolute positioning schemes with simulated GNSS integration. The gray crosses indicate simulated GNSS outages (100-second).
TABLE VI: Horizontal Position RMSEs (m) of Different Positioning Schemes with GNSS RTK integration or with simulated GNSS integration.
Mode
Seq.
GNSS
RTK
VINS
Fusion
DBA
Fusion
M-Fus.
w/o loop
M-Fus.
w/ loop
Real (a)
4.36
2.54
0.78
0.24
0.21
1.46
0.62
0.24
0.13
0.09
Simu. (a)
2.84
3.20
0.69
0.37
9.66
2.94
1.02
0.46
VI-D Reconstruction Results
@@
VII Conclusion
@@ Acknowledgments
This work was supported by the National Science Fund for Distinguished Young Scholars of China (42425401), the National Natural Science Foundation of China (423B240), the National Key Research and Development Program of China (2023YFB3907100), and the China Postdoctoral Science Foundation-Hubei Joint Support Program (2025T041HB). The implemented MASt3R-Fusion is developed by the GREAT Group, School of Geodesy and Geomatics, Wuhan University. The numerical calculations in this paper have been done on the supercomputing system at the Supercomputing Center of Wuhan University.